# Gravitational-Wave Background from Extragalactic Double White Dwarfs for LISA

### v1

@@ Abstract
Results. Our study is in global agreement with previous studies, although we find a lower contribution at high frequencies, due to a different treatment of mass transfer in stellar binaries. We find that the uncertainties in the amplitude are dominated by the star formation model, and to a lesser degree by the binary evolution model. The inclusion of tidal effects and mass transfer episodes in DWDs can change the amplitude of the estimated background up to a factor 3 at the highest frequencies. For all the models we consider, we find that this background could be potentially observable by LISA after 4 years of observations. Under the hypothesis of an homogeneous Universe beyond 200 Mpc, anisotropies associated to the astrophysical population of DWDs will likely not be detectable. We provide phenomenogical fits of the background produced by extragalactic DWDs under different assumptions to be used by the community.
@@ 2.1 LISA sensitivity to stochastic backgrounds
For the rest of the paper we assume the equal correlation hypothesis and the LISA noise model is given by the power-law integrated sensitivity (PI curve, following Thrane & Romano, 2013) for four years of observation time and a fixed value of signal-to-noise ratio SNR=10 (gray solid line on figures (Pol, 2023)). The PI curve is a graphical tool used to represent the sensitivity of GW detectors to stochastic backgrounds characterized by power-law spectra. It indicates the minimum energy density of a stochastic background that a detector can observe across different spectral ranges, assuming the background is isotropic, stationary, and Gaussian.
@@ 2.2 SGWB formalism and computation
where $T_{\mathrm{obs}}=4$ and 10 years is the observation time and $\Omega_{\mathrm{sens}}(f)$ is the effective sensitivity that includes LISA’s instrumental noise the Galactic confusion noise:
@@ 3.1 Predictions for the SGWB from extragalactic DWDs
Figure 1: Spectral energy density of the SGWB from extragalactic DWDs for various population synthesis models (see §2.3) and various SFRD models (see §2.4.1). The different population synthesis models are indicated with the same markers as in Fig. 9. Dotted, solid, and dashed lines correspond to the SFR models from Strolger et al. (2004), Madau & Dickinson (2014), and Madau & Fragos (2017) respectively. The black line represents the LISA sensitivity curve (Colpi et al., 2024), while the gray one shows the PI curve (Thrane & Romano, 2013) for four years of observation time and SNR=10. The light blue curve is the confusion noise from Galactic DWDs (Robson et al., 2019) and the green shaded band is the 90% credible interval of the SGWB from all compact object binaries estimated from observations of GWs (Abbott et al., 2023). As a comparison, the blue shaded band shows the SGWB from extragalactic DWDs estimated in (Farmer & Phinney, 2003) and the pink curve indicates the the best-fit model obtained in (Staelens & Nelemans, 2024). In Appendix, we have discussion and summarize fit of SGWB from extragalactic DWDs of all SFR and stellar model synthesis section B and Table 4.
@@
The spread arising from the different population synthesis models is fairly similar across the different SFRD models and remains consistent with our previous description using the default SFRD model from Madau & Dickinson (2014). It is clear that the choice of SFRD model has a major impact on the global amplitude of the SGWB from extragalactic DWDs. Indeed, as the fraction of stellar binaries that evolve to become DWDs is roughly constant over a large range of metallicities (see Fig. 9), the amplitude of the SGWB from extragalactic DWDs naturally relates to the total mass of stars that have formed in the entire history of the Universe, or the integrated SFR. As the SFRD model from Strolger et al. (2004) predicts notably more star formation at $z>2$ than the two other SFRD models (see the left-hand panel of Fig. 10), this naturally translates into much higher SGWB amplitudes. We find that the use of different SFRD models dominate the uncertainty in the SGWB amplitude, with a factor of up to 6-7, while different models of initial stellar populations and of binary evolution have a more moderate impact, with a factor 2-3 variation in the global amplitude of the predicted SGWB.
@@
These results indicate that the SGWB from extragalactic DWDs could be potentially observed over the course of the LISA mission. However, the degeneracies between the effect of the cosmological SFRD, the properties of initial stellar populations, and the nature of common envelope evolution will make it extremely difficult to constrain any of these physical processes independently. Furthermore, if the SFRD turns out to be larger at high redshifts than the commonly used model of Madau & Dickinson (2014), then the SGWB from extragalactic DWDs could even have a major impact on the sensitivity curve of LISA in the frequency band $[10^{-3},10^{-2}]\ \mathrm{Hz}$.
@@
Stellar Model binfrac1 2.94 4.65
multidim 3.20 5.05
default 2.14 3.39
alpha4 5.54 8.75
@@
Stellar Model binfrac1 2.09 3.31
multidim 2.26 3.58
defaultn 1.52 2.41
alpha4 3.36 5.31
Strolger
Stellar Model binfrac1 13.38 21.15
multidim 13.75 21.74
default 9.63 15.22
alpha4 19.95 31.54
To assess detectability, Table 2 summarizes the computed SNR (Eq 14) for various combinations of stellar evolution models and SFR, considering both 4-year and extended 10-year LISA mission durations. The (Strolger et al., 2004) scenario systematically yields higher SNR, notably reaching $\sim 20$ (alpha4, $T=4$ yr) and further increasing up to $\sim 32$ for a 10-year observation, reinforcing its comparatively strong detection potential. Conversely, scenarios involving the default model paired with the (Madau & Dickinson, 2014) or (Madau & Fragos, 2017) SFR yield substantially lower values ($\sim 1.5-3$ for $T=4$ yr, only modestly improved at $\sim 2-3.5$ over 10 years). These lower SNR values illustrate the substantial observational uncertainties arising from Galactic foreground, instrumental sensitivity limits, and fundamental uncertainties in binary evolution modeling. Therefore, although extending the observation period significantly improves detectability for some model combinations, the detection of the SGWB from extragalactic DWDs by LISA remains conditional upon overcoming these observational and theoretical uncertainties.
@@ 4.1 Uncertainties in modeling the SGWB
The high frequency regime will also be the most affected by tidal torques and/or mass transfer between the DWDs. Following the Toubiana et al. (2024) model, we find up to a factor 2 variation in amplitude at the highest frequencies. While the high frequency regime appears to be the most interesting to study the effects of binary interaction, it is also the least likely to be detected, unless the signal amplitude is higher than predicted or the population of tight systems is more abundant than assumed.
4.2 Detectability
The detectability of the SGWB from extragalactic DWDs by LISA is highly dependent on the combination of stellar evolution and star formation models used. Table 2 shows that the computed SNR varies by over an order of magnitude across all model combinations. For a 4-year mission, values range from $\sim 1.5$ (e.g., default + Madau fragos) up to nearly $20$ (e.g., alpha4 + Strolger), with corresponding increases for a 10-year mission (up to $\sim 32$ in the most favorable case).
This variation reflects a strong sensitivity to assumptions about the common-envelope phase and the global star formation rate. High-SFR models such as Strolger naturally lead to a larger integrated background, while prescriptions such as alpha4, with more efficient CE ejection, increase the number of tight DWDs contributing in the LISA band. In contrast, models based on more conservative assumptions (default or Madau Fragos) systematically yield lower SNR, often remaining below practical detectability limits, even for extended mission durations.
These results confirm that detection is not guaranteed and should not be assumed based on a PI curve. Only a restricted subset of plausible scenarios lead to robust SNR levels. Interpreting any potential detection will require careful modelling of foregrounds and inclusion of the extragalactic DWD background in future data analysis pipelines. A more detailed analysis of its detectability and the associated implications will be the subject of future work.
@@ 5 Conclusions
  • The signal is detectable by LISA within 4 years in optimistic scenarios (e.g., Strolger + alpha4), with SNRs exceeding 20 and rising above 30 for 10-year integrations. For most of the configurations we explored, the signal will remain well below the nominal detectability threshold.
@@
Reducing model uncertainty requires improved treatment of binary evolution, particularly during the common-envelope phase. Multi-messenger comparisons, such as those discussed in van Zeist et al. (2025), offer promising avenues to calibrate population synthesis outputs using electromagnetic constraints. Further cross-validation between independent codes (e.g., BPASS (Stanway & Eldridge, 2023; Byrne et al., 2025)) will also help identify robust features in SGWB predictions.
@@ A.1 Physical properties of DWDs at formation and formation efficiency
Figure 10: Left: SFRD in the Universe as a function of redshift. Different colours indicate different estimates based on different assumptions on the initial mass function of stellar populations or on extinction correction (Strolger et al., 2004; Madau & Dickinson, 2014; Madau & Fragos, 2017). Right: Metallicity-specific SFRD as a function of redshift for different values of metallicity, computed as the combination of the SFRD from Madau & Dickinson (2014) with the metallicity distribution of Neijssel et al. (2019). The dashed, dot-dashed, and dotted lines show the metallicity-specific SFRD around metallicity $\mathrm{Z}_{\odot}$, $0.1\times\mathrm{Z}_{\odot}$, and $0.01\times\mathrm{Z}_{\odot}$ respectively.
@@ Appendix B Estimating the SGWB Using a Smoothly Broken Power Law fit
where $A$ is the amplitude, $f_{\mathrm{b}}$ is the frequency location of the knee, $\alpha_{1}$ and $\alpha_{2}$ are the spectral indices before and after this frequency, and $\delta$ controls the smoothness of the transition. The implementation of this model was facilitated using the smoothly broken power law from the astropy library (Robitaille et al., 2013; Price-Whelan et al., 2022).
In our analysis, the spectral fitting of the SGWB from extragalactic DWDs is carried out using the dynesty library (Speagle, 2020), which implements a nested-sampling Markov chain Monte Carlo (MCMC) algorithm. This Bayesian approach yields posterior distributions for the model parameters, their uncertainties, and the model evidence for possible comparison. The likelihood is defined as the sum of squared differences between the model and the simulated SGWB from extragalactic DWDs spectrum, weighted by a standard deviation that is also estimated during inference.
This methodology follows closely the approach developed in previous studies of the SGWB for LISA, which employed similar MCMC-based strategies for spectral parameter estimation (Boileau et al., 2021, 2022). In particular, Boileau et al. (2021) introduced an adaptive MCMC algorithm to analyze simulated LISA data, demonstrating the effectiveness of Bayesian inference in separating astrophysical and cosmological background components. Building on this, Boileau et al. (2022) extended the Bayesian MCMC framework to assess the detectability of a cosmologically-generated background in the presence of a foreground, highlighting the relevance of such techniques for SGWB analyses. By adopting a nested-sampling MCMC approach in this work, we ensure methodological continuity with these studies, which strengthens the reliability of our SGWB from extragalactic DWDs spectral estimation and enables direct comparison with previous results in the LISA context.
@@
Acknowledgements.
We sincerely thank the group of Gijs Nelemans for their valuable assistance and the constructive discussions that helped us better understand the differences between our results. Their insights were instrumental in refining our analysis and interpretations. G.B. thanks the Centre national d’études spatiales (CNES) for support for this research. This project has received financial support from the CNRS through the MITI interdisciplinary programs through its exploratory research program. A.L., T.B. and N.C. acknowledge support by the French Agence Nationale de la Recherche. T.B. and A.T. are supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. A.T. is supported by MUR Young Researchers Grant No. SOE2024-0000125.

### v2

@@ Abstract
Results. Our study is in global agreement with previous studies, although we find a lower contribution at high frequencies, due to a different treatment of mass transfer in stellar binaries. We find that the uncertainties in the amplitude are dominated by the star formation model, and to a lesser degree by the binary evolution model. The inclusion of tidal effects and mass transfer episodes in DWDs can change the amplitude of the estimated background up to a factor 3 at the highest frequencies. For all the models we consider, we find that this background would be easily detectable, with SNR values from 100 to more than 1000 by LISA after 4 years of observations. Under the hypothesis of an homogeneous Universe beyond 200 Mpc, anisotropies associated to the astrophysical population of DWDs will likely not be detectable. We provide phenomenogical fits of the background produced by extragalactic DWDs under different assumptions to be used by the community.
@@ 2.1 LISA sensitivity to stochastic backgrounds
For the rest of the paper we assume the equal correlation hypothesis and the LISA noise model is given by the power-law integrated sensitivity (PI curve, following Thrane & Romano, 2013) for four years of observation time and a fixed value of signal-to-noise ratio SNR=10 (gray solid line on figures, see e.g. Babak et al. (2021)). The PI curve is a graphical tool used to represent the sensitivity of GW detectors to stochastic backgrounds characterized by power-law spectra. It indicates the minimum energy density of a stochastic background that a detector can observe across different spectral ranges, assuming the background is isotropic, stationary, and Gaussian.
@@ 2.2 SGWB formalism and computation
where $T_{\mathrm{obs}}=4$ and 10 years is the observation time and $\Omega_{\mathrm{sens}}(f)$ is the effective sensitivity that includes LISA’s instrumental noise and the Galactic confusion noise:
@@ 3.1 Predictions for the SGWB from extragalactic DWDs
Figure 1: Spectral energy density of the SGWB from extragalactic DWDs for various population synthesis models (see §2.3) and various SFRD models (see §2.4.1). The different population synthesis models are indicated with the same markers as in Fig. 9. Dotted, solid, and dashed lines correspond to the SFR models from Strolger et al. (2004), Madau & Dickinson (2014), and Madau & Fragos (2017) respectively. The black line represents the LISA sensitivity curve (Colpi et al., 2024), while the gray one shows the PI curve (Thrane & Romano, 2013) for four years of observation time and SNR=10. The light blue curve is the confusion noise from Galactic DWDs (Robson et al., 2019) and the green shaded band is the 90% credible interval of the SGWB from all compact object binaries estimated from observations of GWs (Abbott et al., 2023). As a comparison, the blue shaded band shows the SGWB from extragalactic DWDs estimated in (Farmer & Phinney, 2003) and the pink curve indicates the the best-fit model obtained in (Staelens & Nelemans, 2024). In Appendix B, we discuss and summarize the fits of SGWB from extragalactic DWDs of all SFR models and stellar population synthesis models as presented in Table 4.
@@
The spread arising from the different population synthesis models is fairly similar across the different SFRD models and remains consistent with our previous description using the default SFRD model from Madau & Dickinson (2014). It is clear that the choice of SFRD model has a major impact on the global amplitude of the SGWB from extragalactic DWDs. Indeed, as the fraction of stellar binaries that evolve to become DWDs is roughly constant over a large range of metallicities (see Fig. 9), the amplitude of the SGWB from extragalactic DWDs naturally relates to the total mass of stars that have formed in the entire history of the Universe, or the integrated SFR. As the SFRD model from Strolger et al. (2004) predicts notably more star formation at $z>2$ than the two other SFRD models (see the left-hand panel of Fig. 10), this naturally translates into much higher SGWB amplitudes. We find that the use of different SFRD models dominate the uncertainty in the SGWB amplitude, with a factor of up to 6-7, while different models of initial stellar populations and of binary evolution have a more moderate impact, with a factor 2-3 variation in the global amplitude of the predicted SGWB. This variation highlights the strong sensitivity of the SGWB from extragalactic DWDs predictions to assumptions about the common-envelope phase and the global star formation rate.
@@
Strolger
Stellar Model default 748 1182
binfrac1 1037 1640
multidim 1076 1701
alpha4 1493 2361
@@
Stellar Model default 155 244
binfrac1 212 335
multidim 233 369
alpha4 386 610
@@
Stellar Model default 111 176
binfrac1 152 241
multidim 167 264
alpha4 279 441
Table 2 shows that the computed SNR varies by more than an order of magnitude across models, from $\sim 100$ for conservative scenarios (e.g., default + Madau Fragos) up to nearly $\sim 1500$ (e.g., alpha4 + Strolger) for a 4-year mission, and reaching $\sim 2400$ for a 10-year observation in the most favorable case.
Even scenarios involving the default model paired with the Madau & Dickinson (2014) or Madau & Fragos (2017) SFR yield SNR values well above unity ranging from $\sim 100$ to $\sim 400$ for $T=4$ yr and up to $\sim 600$ for a 10-year mission. These very high SNR values demonstrate that , in all scenarios, the detection of the SGWB from extragalactic DWDs by LISA is virtually guaranteed (see Table 2), regardless of Galactic foreground or instrumental sensitivity limits. However, the degeneracies between the effect of the cosmological SFRD, the properties of initial stellar populations, and the nature of common envelope evolution will remain challenging to disentangle, as detection alone will not suffice to constrain these physical processes independently. These fundamental uncertainties in binary evolution modeling and star formation histories will affect the interpretation and parameter inference of the observed background. Therefore, while extending the observation period further increases SNR, the focus should shift toward accurately modeling these astrophysical uncertainties to fully exploit the detection, notably the position of the typical frequency at which the signal starts to drop, referred to here as the knee, which will have an impact on the astrophysical models. Furthermore, if the SFRD turns out to be larger at high redshifts than the commonly used model of Madau & Dickinson (2014), then the SGWB from extragalactic DWDs could have an even stronger impact on the effective sensitivity curve of LISA, particularly in the frequency band $[10^{-3},10^{-2}]\ \mathrm{Hz}$.
@@ 4.1 Uncertainties in modeling the SGWB
The high frequency regime will also be the most affected by tidal torques and/or mass transfer between the DWDs. Following the Toubiana et al. (2024) model, we find up to a factor 2 variation in amplitude at the highest frequencies. While the high frequency regime appears to be the most interesting to study the effects of binary interaction, it is also the least likely to be detected, unless the signal amplitude is higher than predicted or the population of tight systems is more abundant than assumed. Given the high SNRs predicted, it may possible to detect signatures of tidal interactions or mass transfer effects in the observed signal.
4.2 Implications of the knee frequency
Beyond the overall detectability of the SGWB from extragalactic DWDs, the knee frequency provides particularly valuable astrophysical information. This knee frequency is primarily determined by the existence of very massive DWDs with extremely low orbital separations. In addition to the natural scaling of orbital frequency with the mass of the binary components, white dwarfs have the particularity of seeing their radius decrease as their mass increases (see e.g. Kawaler, 1997). This implies that more massive DWDs can reach smaller orbital separations before one starts filling its Roche lobe, hence increasing again the maximum frequency at which they can emit GWs before their merger. To contribute to the high frequency part of the energy density power spectrum, these binaries must also exist at low redsfhits (otherwise their received frequency would be shifted to smaller values). The location of the frequency knee in the SGWB from extragalactic DWDs is therefore based on the possibility and efficiency of binary stellar evolution to produce two massive WDs very close to each other.
Models with more efficient envelope ejection (higher $\alpha$ or $\lambda$ parameters) tend to produce wider post-CE separations, but also allow some binaries to survive the CE evolution phase, whereas they would merge during this phase in the case of less efficient energy transfer to the envelope. This explains why the SGWB from extragalactic DWDs predicted with our alpha4 COSMIC model extends to slightly higher frequencies than our default model (see Fig. 1 and Table 4).
This feature is explored in our comparison with the population synthesis code SeBa in the previous subsection (see e.g. the model $\lambda_{\mathrm{CE}}=0.5\ \&\ \tilde{q}_{\mathrm{crit}}$ and the results of Staelens & Nelemans (2024) presented in Fig. 6).
Therefore, the precise location of this spectral break encodes direct constraints on CE physics, which remains one of the largest uncertainties in binary stellar evolution. Observing the knee frequency with LISA would thus offer an opportunity to constrain the physical prescriptions used for CE evolution and improve the predictions of population synthesis models. Future analyses should therefore focus on precise modelling and measurement of the knee frequency, together with the overall amplitude, to enhance our understanding of binary evolution physics using the SGWB from extragalactic DWDs observed by LISA.
@@ 5 Conclusions
  • The signal is detectable by LISA across all scenarios explored, with SNRs ranging from $\sim 100$ to $\sim 400$ for conservative models ans exceeding $\sim 1000$ for optimistic models (e.g., Strolger + alpha4). For 10-year integration, the SNR can reach up to $\sim 2400$, indicating that the extragalatic DWD background will be robustly observed by LISA under all astrophysical assumption considered here.
@@
Reducing model uncertainty requires improved treatment of binary evolution, particularly during the common-envelope phase. Multi-messenger comparisons, such as those discussed in van Zeist et al. (2025), offer promising avenues to calibrate population synthesis outputs using electromagnetic constraints. Further cross-validation between independent codes (e.g., BPASS (Stanway & Eldridge, 2023; Byrne et al., 2025)) will also help identify robust features in SGWB predictions. Interpreting this detection will require careful modelling of foregrounds and explicit inclusion of the SGWB from extragalactic DWDs in future LISA data analysis pipelines. A more detailed analysis of its detectability and implications will be the subject of future works.
@@
Acknowledgements.
We sincerely thank the group of Gijs Nelemans for their valuable assistance and the constructive discussions that helped us better understand the differences between our results. We would also like to thank Arianna Renzini and Riccardo Buscicchio for their priceless comments and advice. Their insights were instrumental in refining our analysis and interpretations. G.B. thanks the Centre national d’études spatiales (CNES) for support for this research. This project has received financial support from the CNRS through the MITI interdisciplinary programs through its exploratory research program. A.L., T.B. and N.C. acknowledge support by the French Agence Nationale de la Recherche. T.B. and A.T. are supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. A.T. is supported by MUR Young Researchers Grant No. SOE2024-0000125.
@@ A.1 Physical properties of DWDs at formation and formation efficiency
Figure 10: Left: SFRD in the Universe as a function of redshift. Different colours indicate different estimates based on different assumptions on the initial mass function of stellar populations or on extinction correction (Strolger et al. 2004; Madau & Dickinson 2014; Madau & Fragos 2017). Right: Metallicity-specific SFRD as a function of redshift for different values of metallicity, computed as the combination of the SFRD from Madau & Dickinson (2014) with the metallicity distribution of Neijssel et al. (2019). The dashed, dot-dashed, and dotted lines show the metallicity-specific SFRD around metallicity $\mathrm{Z}_{\odot}$, $0.1\times\mathrm{Z}_{\odot}$, and $0.01\times\mathrm{Z}_{\odot}$ respectively.
@@ Appendix B Estimating the SGWB Using a Smoothly Broken Power Law fit
where $A$ is the amplitude, $f_{\mathrm{b}}$ is the frequency location of the knee, $\alpha_{1}$ and $\alpha_{2}$ are the spectral indices before and after this frequency, and $\delta$ controls the smoothness of the transition. The implementation of this model was facilitated using the smoothly broken power law from the astropy library (Robitaille et al. 2013; Price-Whelan et al. 2022).
In our analysis, the spectral fitting of the SGWB from extragalactic DWDs is carried out using the dynesty library (Speagle 2020), which implements a nested-sampling Markov chain Monte Carlo (MCMC) algorithm. This Bayesian approach yields posterior distributions for the model parameters, their uncertainties, and the model evidence for possible comparison. The likelihood is defined as the sum of squared differences between the model and the simulated SGWB from extragalactic DWDs spectrum, weighted by the standard deviation of data accross the fitted region.
This methodology follows closely the approach developed in previous studies of the SGWB for LISA, which employed similar MCMC-based strategies for spectral parameter estimation (Boileau et al. 2021, 2022). In particular, Boileau et al. (2021) introduced an adaptive MCMC algorithm to analyze simulated LISA data, demonstrating the effectiveness of Bayesian inference in separating astrophysical and cosmological background components. Building on this, Boileau et al. (2022) extended the Bayesian MCMC framework to assess the detectability of a cosmologically-generated background in the presence of a foreground, highlighting the relevance of such techniques for SGWB analyses. By adopting a nested-sampling MCMC approach in this work, we ensure methodological continuity with these studies, which strengthens the reliability of our SGWB from extragalactic DWDs spectral estimation and enables direct comparison with previous results in the LISA context.